OpenStudy (anonymous):

Percent proportion problem. 6.25% of 190 is is what number? I don't just want an answer. I need step by step directions so I can figure it out on my own. But I need it dummied down for me.

OpenStudy (amistre64):

1% means, 1 out of 100 so if we divide 190 into 100 parts, what is the value of 1 part?

OpenStudy (amistre64):

when know the value of 1 part, all we do is add up however many of those we need. and when we add the same thing multiple times ... its called multiplication.

OpenStudy (texaschic101):

there is a percentage formula that works for every percent problem. is/of = %/100 6.25% of 190 is what number ? of = 190 is = x % = 6.25 now we sub x/190 = 6.25/100 --- this is now a proportion, so we will cross multiply 100x = (190)(6.25) 100x = 1187.50 x = 1187.50/100 x = 11.875 check.. 6.25% of 190 = .0625(190) = 11.875 and this formula works for every percent problem....just be sure you sub in the numbers in the right spots

OpenStudy (anonymous):

Awesome! Thanks so much! You've all been a lot of help. But especially you, Texaschic.

OpenStudy (texaschic101):

glad to help :)